Sugar syrup of a mass 214.2g contain 34.2g of sugar [C12 H22 O11], Calculate molarity and mole fraction.
 

Grade:

2 Answers

Ashish
5 years ago
Weight of sugar syrup is = 214.2 g
Weight of sugar in syrup = 34.2 g
Weight of water in sugar syrup = 214.2-34.2 =180g
Molecular weight of Sugar =342g
Mole fraction of sugar=
 
 
34.2/342(180+18)+34.2/342(180+18)+34.2/342
Ashish
5 years ago
In previous answer (yah answer chut gaya tha)
Molecular weight of Sugar= 342g
Molarity=
34.2*1000/342*180= 0.56